OpenStudy (ladylauren):

Freddy reads an equal number of pages of a book every week. The graph below shows the number of pages of the book left to read, y, after x weeks:

OpenStudy (chrishogan):

y = –5x + 200

OpenStudy (ladylauren):

Which equation best models the relationship between x and y?

OpenStudy (ladylauren):

ya

OpenStudy (chrishogan):

Your answer should be y = –5x + 200 IS that one of your answer choices?

OpenStudy (ladylauren):

yes

OpenStudy (chrishogan):

Aight that's the answer. Need help with any other problems?

OpenStudy (ladylauren):

yes can you help

OpenStudy (chrishogan):

Of course.

OpenStudy (ladylauren):

The graph shows the price, in dollars, of different numbers of sweet breads at Alan's store. The table shows the price, in dollars, of different numbers of shortcakes at the same store.

OpenStudy (ladylauren):

Shortcake Number of Shortcake Price of Shortcake 5 45 10 90 15 135 20 180

OpenStudy (ladylauren):

How many dollars more is the price of a shortcake than the price of a sweet bread at David's store?

OpenStudy (ladylauren):

$4 $5 $9 $25

OpenStudy (chrishogan):

$25 should be your answer.

OpenStudy (ladylauren):

can you help me with 1 more?

OpenStudy (chrishogan):

Of course.

OpenStudy (chrishogan):

Start a new question though.

OpenStudy (ladylauren):

Thomas wants to rent a lawn mower. He has to pay a fixed base cost plus a daily rate for renting the lawn mower. The table shows the amount of money, y, in dollars, that Thomas has to pay for renting the lawn mower for x days: Lawn Mower Rental Number of days Rent (dollars) x y 0 12 1 21 2 30 3 39 4 48 Which equation best shows the relationship between x and y? y = x + 21 y = 9x + 21 y = 9x + 12 y = x + 12

OpenStudy (chrishogan):

your answer would be y=9x+12

OpenStudy (chrishogan):

So, C.

OpenStudy (ladylauren):

ok thx so much

OpenStudy (chrishogan):

Np. Just do @chrishogan on one of your questions if you ever need help again.

OpenStudy (ladylauren):

ok thx again